2. Ammonia gas reacts with sulfuric acid to form the important fertilizer, ammonium sulfate. What mass of ammonium sulfate can be produced from 45.0 L of ammonia at 10 °C and 110 kPa?

3. Z-dawg finds that 0.06g of magnesium in excess hydrochloric acid produces hydrogen gas.
The lab is at 22.1C and 100.8kpa.
A. What is the balanced reaction?
B. What is the volume of hydrogen gas produced?
C. Z-dwag collects 71.90mL of gas. What is the percent yield of this experiment?
D. What are two errors that could have occured in this lab that lead to this percent error?

4. A sample of zinc was placed in 100.0 mL of 1.50 mol/L hydrochloric acid and reacted all of
the acid. What was the mass of the zinc sample?

5. What is the concentration of a 20.0mL solution of hydrofluoric acid that reacts with 84.5mL of
a 0.400mol/L solution of barium hydroxide?
6. What is the volume of hydrogen made at 21C and 110kpa from a 30.0mL, 0.200mol/L
solution of hydrobromic acid reacting with excess aluminium
1
Expert's answer
2020-05-27T13:34:17-0400

2) "pV=nRT=>\nn=pV\/RT=2.104 mol=>n((NH_4)_2SO_4)=n\/2=1.052mol=>"

"=>m((NH_4)_2SO_4)=n((NH_4)_2SO_4)*M((NH_4)_2SO_4)=1.052*132g=138.86g"


3)

A) "Mg+2HCl=MgCl_2+H_2"

B) "n=0.06\/24=0.0025mol=n(H_2)=>V=nRT\/p=0.00006088m^3=60.88mL"

C) "yield: 71.90\/60.88=1.1810 (118.10 ^o\/_o)"

D) Thermometer and barometer are broken


4) "Zn+2HCl=ZnCl_2+H_2\n,n(HCl)=CV=1.5*0.1mol=0.15mol"

"n(Zn)=n(HCl)\/2=0.075mol=>m(Zn)=n(Zn)*M(Zn)=0.075*65.38g=4.9g"


5) "Ba(OH)_2+2HF=BaF_2+2H_2O"

"n(HF)=2*n(Ba(OH)_2)= 2*0.0845L* 0.400mol\/L =0.0676mol="

"=C(HF)*V(HF)=>C(HF)=n(HF)\/V(HF)=0.0676\/0.02=3.38M"


6) "2Al+6HBr=2AlBr_3+3H_2"

"n(HBr)=0.0300L*0.200mol\/L=0.006 mol=>n(H_2)=3*0.006\/6=0.003 mol"

"V=nRT\/p=66.7mL"